OpenStudy (anonymous):

2 opposite current i=I[sin(wt)] passed through a resistor..find the thermal energy developed

OpenStudy (anonymous):

where are the two currents ??

OpenStudy (anonymous):

the two currents are produced by 2 ac source..one is i=Isinwt another i=-Isinwt they pass through a resistor R ..find the heat produced

OpenStudy (anonymous):

i am not sure as current add up like vectors and result in no current and hence no heat or thermal energy add up like scalars !

OpenStudy (anonymous):

@srijit What you are given is that you have an ac current source. Of course, the heat will be produced. Effect of AC current source will be that after every half cycle, the direction of current will be changed. Heat will be produced during every half cycle. And as you know very well due to passage of current, heat will be produced. This is what I think If you still have any doubt, let me know.

OpenStudy (anonymous):

And current don't add up like vectors.

OpenStudy (radar):

If any heat is produced, it would be from the resultant of those two currents. What is the resultant current?

OpenStudy (anonymous):

@radar resultant current is Isinwt Am i wrong??

OpenStudy (radar):

It would depend on t ( the phase of the currents). The problem needs to be worded so it is clear. If the two currents are equal in magnitude and exactly 180 degrees out of phase, then the resultant would be 0.

OpenStudy (anonymous):

the currents are out of phase as i said -one Isinwt and other -Isinwt and resultant 0..

OpenStudy (theeric):

I'm not very good with electronics yet, but I see things from srijit's initial point of view. I think if currents oppose one another (which seems like a contradiction in itself) then they would cancel out. Like addition. \(I\sin(\omega t)+ \left(-I\sin(\omega t)\right)=I\sin(\omega t)-I\sin(\omega t)=0\) for all \(t\) So, no current?

OpenStudy (anonymous):